For the random variable X with the probability function below, answer the following questions.

f(x)= { 3x^2 , 0<x<1

{ 0, otherwise

a. Find the cdf of X.

b. Find the probability that X is less than 0.5.

c. Find E(X).


1
Expert's answer
2021-01-26T04:32:35-0500

The probability function "f(x)=\begin{cases}
3x^2,\ \ \ 0<x<1
\\0, \ \ \text{otherwise}
\end{cases}"

a. The CDF of X: "F(x)=P(X\leq x)=\int\limits _{-\infin}^x f(t)dt"

If "x\leq 0:" "F(x)= \int\limits _{-\infin}^x 0dt =0"

If "0<x<1" : "F(x)= \int\limits _{0}^x 3t^2dt =t^3\big|^x_0=x^3"

If "1\leq x" : "F(x)= \int\limits _{0}^1 3t^2dt =t^3\big|^1_0=1"

"F(x)=\begin{cases}
0,\ \ x\leq 0
\\
x^3,\ \ 0<x<1
\\
1,\ \ 1\leq x
\end{cases}"


b. "P(X<0.5)= F(0.5)=(0.5)^3=0.125"

c. "E(x)=\int\limits _{-\infin}^{+\infin}xf(x)dx=\int\limits _0^1 3x^3dx=\frac{3}{4}x^4\big|^1_0=\frac{3}{4}"


Answer:

a. "F(x)=\begin{cases}
0,\ \ x\leq 0
\\
x^3,\ \ 0<x<1
\\
1,\ \ 1\leq x
\end{cases}"

b. "P(X<0.5)=0.125"

c. "E(x)=\frac{3}{4}"